OpenStudy (vcabral1134):

Here are the answers to my test Which is the direct linear variation equation for the relationship? y varies directly with x and y = 12 when x = 4. A.y = x – 8 B.y = 3x C.y = 2x + 4 D.y = x + 8 B is the correct answer

OpenStudy (vcabral1134):

Which is the quadratic variation equation for the relationship? y varies directly with x2 and y = 36 when x = 3. A. y = 12x B. y = 1/4 x^2 C. y = x^2 + 25 D. y = 4x^2 D is the correct answer

OpenStudy (adajiamcneal):

@amistre64 coc

OpenStudy (amorfide):

correct

OpenStudy (vcabral1134):

According to the two-way table, how many people are in Group 2? Category A Category B Group 1 24 17 Group 2 19 23 The answer is 42

OpenStudy (amorfide):

correfct

OpenStudy (anonymous):

Both of them were correct, tell Brianna i say hi.

OpenStudy (vcabral1134):

The two-way table shows the number of men and women who voted for Candidate A or B. What percent of people who voted for Candidate A were women? Express your answer to the nearest whole percent. Men Women Candidate A 1253 1921 Candidate B 1564 1723 61%

OpenStudy (vcabral1134):

ok @StudyGuidance

OpenStudy (anonymous):

The two-way table shows the number of men and women who voted for Candidate A or B. What percent of people who voted for Candidate A were women? Express your answer to the nearest whole percent. Men Women Candidate A 1253 1921 Candidate B 1564 1723 61% you had that one correct let me check the mean of the next one

OpenStudy (vcabral1134):

What is the range of Variable A for the cluster? 1 to 4 http://static.k12.com/calms_media/media/1549000_1549500/1549023/1/3f26799d6c095052e3eb7012278b03a602b5946c/MS_IMC-141015-108156.jpg

OpenStudy (anonymous):

The mean is 4.4, yes

OpenStudy (vcabral1134):

A company tracked the number of job-related injuries each month. The scatter plot shows the data. What is the month for the outlier? http://static.k12.com/calms_media/media/1549000_1549500/1549025/2/6a631703c74709107e3c4b3c0639279fef8b5130/MS_IMC-141015-108158.jpg The answer is 3

OpenStudy (anonymous):

i got 1 - 6 on the range so im not sure

OpenStudy (vcabral1134):

@StudyGuidance I already took the test I am just putting the answer that are correct

OpenStudy (anonymous):

oh ok so im wrong on that one sorry :)

OpenStudy (vcabral1134):

Which associations best describe the data in the table? Variable X 1 2 5 6 3 4 6 7 5 Variable Y 120 115 89 40 100 80 70 50 60 Choose exactly two answers that are correct. A. linear association B. negative association C. nonlinear association D. positive association A and B

OpenStudy (vcabral1134):

he scatter plot shows the linear model for Variable A and Variable B. What is a possible value for Variable A when Variable B is 3? http://static.k12.com/calms_media/media/1549000_1549500/1549029/1/20e9d3aae95e260a141a8a44aa1fd3b39b9703e4/MS_IMC-141015-1081512.jpg The answer is 8

OpenStudy (vcabral1134):

A service organization held a fundraiser to increase the amount of funds it had to use. The scatter plot shows the amount of money raised over time. According to the model, how much money did the service organization begin with? http://static.k12.com/calms_media/media/1549000_1549500/1549030/1/1ca46b7c908803ff997c4518080ef2329748495f/MS_IMC-141015-1081513.jpg The answer is $200

OpenStudy (vcabral1134):

@StudyGuidance Bri said Hi back

OpenStudy (anonymous):

cool thx!